# A Coincidence Algebra Bundle for Decay Quivers: An Algebraic Approach to Gamma-ray Spectroscopy

### v1

@@ 3.2 Coincidence Algebra Bundle
The transitions of interest are given in the center term where summing in is accounted for with potential summing out between disconnected paths. Each of these paths are connected to their possible branching levels above, with summing out probabilities connecting them, and connected to the ground state, also through summing out probabilities.
Example 3.1.
Let $h_{ij}=\varphi_{h}(x_{ij})$ (detected transition) and $o_{ij}=\varphi_{o}(x_{ij})$ (summing out transition). The probability vector with detection maps for the 4 level decay quiver is given by the following:
$\Gamma(d)$ $=$ $\tilde{b}\cdot_{\varphi_{o}(\tau)}\overline{\tilde{\varphi}_{h}(\tau)}^{n_{\varphi_{o}(\tau)}}\cdot_{\varphi_{o}(\tau)}\widehat{\epsilon_{0}}$
$=$ $\color[rgb]{1,0,0}{f_{3}h_{30}\widehat{p}_{30}+f_{3}h_{31}o_{10}\widehat{p}_{31}\otimes\widehat{\epsilon}_{0}+f_{3}h_{32}o_{20}\widehat{p}_{32}\otimes\widehat{\epsilon}_{0}+f_{3}o_{32}h_{21}o_{10}\widehat{\epsilon}_{3}\otimes\widehat{p}_{21}\otimes\widehat{\epsilon}_{0}+}$
$\color[rgb]{1,0,0}{f_{3}o_{32}h_{21}o_{10}\widehat{\epsilon}_{3}\otimes\widehat{p}_{21}\otimes\widehat{\epsilon}_{0}+f_{3}o_{32}h_{20}\widehat{\epsilon}_{3}\otimes\widehat{p}_{20}+f_{2}h_{20}\widehat{p}_{20}}$
$\color[rgb]{1,0,0}{+f_{3}(o_{31}+o_{32}o_{21})h_{10}\widehat{\epsilon}_{3}\otimes\widehat{p}_{10}+f_{2}o_{32}o_{21}h_{10}\widehat{\epsilon}_{2}\otimes\widehat{p}_{10}+f_{1}h_{10}\widehat{p}_{10}}$
$\color[rgb]{0,0,1}{f_{3}h_{31}h_{10}\widehat{p_{31}p_{10}}+f_{3}h_{32}h_{20}\widehat{p_{32}p_{20}}+f_{3}h_{32}o_{21}h_{10}\widehat{p}_{32}\otimes\widehat{p}_{10}+}$
$\color[rgb]{0,0,1}{f_{3}h_{32}h_{21}o_{10}\widehat{p_{32}p_{21}}\otimes\widehat{\epsilon}_{0}+f_{3}o_{32}h_{21}h_{10}\widehat{\epsilon}_{3}\otimes\widehat{p_{21}p_{10}}+f_{2}h_{21}h_{10}\widehat{p_{21}p_{10}}}$
$f_{3}h_{32}h_{21}h_{10}\widehat{p_{32}p_{31}p_{10}}$.
The red terms correspond to all events where one gamma in the decay is detected at full energy with summing out probabilities accounted for and different branching terms are separated. The blue terms give the events where two gammas sum together with full energy and summing out of the remaining gammas is accounted for is applicable. Finally, the black term gives the triple summing event.
4 Conclusion
@@
Motivated by the need to extend the generality of the transition matrix formalism traditionally used to model nuclear decay schemes, We have demonstrated how the path algebra of a decay quiver gives the underlying algebra for transitions matrices. This allowed for an extension into a tensor decay space which we have restricted by defining a coincidence vector space upon which we build the coincidence algebra. The coincidence algebra allows for a more general algebra for calculating coincidence probabilities as it allows for multiplication between non connected paths in the decay quiver and in addition separates different branching terms. With the path algebra being the space with which nuclear levels schemes live, we have define an algebra bundle over the path algebra where each decay vector defines the structure constants on the coincidence algebra at the fibre. We have showed, at the very least that this algebra allows for additional coincidence probability terms to be calculated. We propose, however, that this method of taking fibre bundle of quivers, or potentially even quiver representations, may be a fruitful method for modeling at nuclear levels schemes. This formalism may potentially be extended to include angular correlations as loops in the base space and potentially and energy and spin representations to the structure. This formalism may also have an interesting intersection with geometric nuclear theory.
Acknowledgements
This work was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC).

### v2

@@ III.2 Coincidence Algebra Bundle
The transitions of interest are given in the center term where summing in is accounted for with potential summing out between disconnected paths. Each path is connected to their possible branching levels above and to the ground state below through the summing out probabilities connecting between. As an example, the probability vector with detection maps for the 4 level decay quiver given by equation 42 is shown in Table 1. The basis vectors are separated by the number of fully energy gammas they represent. The coefficients are shown in the columns and are distinguished by the branch they decayed from.
Basis Vector $f_{3}$ $f_{2}$ $f_{1}$
$\widehat{\epsilon}_{3}\otimes\widehat{p}_{30}$ $f_{3}h_{30}$
$\widehat{\epsilon}_{3}\otimes\widehat{p}_{31}\otimes\widehat{\epsilon}_{0}$ $f_{3}h_{31}o_{10}$
$\widehat{\epsilon}_{3}\otimes\widehat{p}_{32}\otimes\widehat{\epsilon}_{0}$ $f_{3}h_{32}o_{20}$
$\widehat{\epsilon}_{3}\otimes\widehat{p}_{21}\otimes\widehat{\epsilon}_{0}$ $f_{3}o_{32}h_{21}o_{10}$
$\widehat{\epsilon}_{3}\otimes\widehat{p}_{20}$ $f_{3}o_{32}h_{20}$
$\widehat{\epsilon}_{3}\otimes\widehat{p}_{10}$ $f_{3}(o_{31}+o_{32}o_{21})h_{10}$
$\widehat{\epsilon}_{2}\otimes\widehat{p}_{20}$ $f_{2}h_{20}$
$\widehat{\epsilon}_{2}\otimes\widehat{p}_{10}$ $o_{21}h_{10}$
$\widehat{\epsilon}_{1}\otimes\widehat{p}_{10}$ $f_{1}h_{10}$
$\widehat{\epsilon}_{3}\otimes\widehat{p_{31}}\otimes\widehat{p_{10}}$ $f_{3}h_{31}h_{10}/N$
$\widehat{\epsilon}_{3}\otimes\widehat{p_{32}}\otimes\widehat{p_{20}}$ $f_{3}h_{32}h_{20}/N$
$\widehat{\epsilon}_{3}\otimes\widehat{p}_{32}\otimes\widehat{p}_{10}$ $f_{3}h_{32}o_{21}h_{10}/N$
$\widehat{\epsilon}_{3}\otimes\widehat{p_{32}}\otimes\widehat{p_{21}}\otimes\widehat{\epsilon}_{0}$ $f_{3}h_{32}h_{21}o_{10}/N$
$\widehat{\epsilon}_{3}\otimes\widehat{p_{21}}\otimes\widehat{p_{10}}$ $f_{3}o_{32}h_{21}h_{10}/N$
$\widehat{\epsilon}_{2}\otimes\widehat{p_{21}}\otimes\widehat{p_{10}}$ $f_{2}h_{21}h_{10}/N$
$\widehat{\epsilon}_{3}\otimes\widehat{p_{32}}\otimes\widehat{p_{21}}\otimes\widehat{p_{10}}$ $f_{3}h_{32}h_{21}h_{10}/N^{2}$
Table 3: Probability vector $\Gamma(d)$ coefficients for a 4 level decay quiver with summing out probabilities $o_{ij}=\varphi_{o}(x_{ij})$ and detected transition $h_{ij}=\varphi_{h}(x_{ij})$. The basis vectors are separated by the branch they decayed from. The first block contains the basis vectors where 1 gamma is detected at full energy; the second block contains the basis vectors where 2 gammas are detected at full energy; the third block contains the only basis vector where 3 gammas are detected at full energy.
To calculate the detected coincidence probability vector $\Gamma_{C}(d)$ with summing corrections, we must first define a new detection map that allows for the second gamma (with the first being the gate) to be detected–that is, it does not sum with the first gamma. To do this we continue to use the isotropic approximation and say the probability that the second gamma does not sum with the first is given by (N1)/N(N-1)/N. We therefore define the gated hit detection map as
$\varphi_{g}=\frac{\varphi_{h}(N-1)}{N}.$ (44)
Again, with $N$ being the number of detectors. The detected coincidence probability vectors is therefore
$\Gamma_{C}(d)=\tilde{b}\cdot\overline{\tilde{\varphi}_{h}(\tau)}^{n_{\varphi_{o}(\tau)}}\cdot\overline{\tilde{\varphi}_{g}(\tau)}^{n_{\varphi_{o}(\tau)}}\cdot\widehat{\epsilon_{0}}$, (45)
where the multiplication $\cdot:=\cdot_{\varphi_{o}(\tau)}$ is over the summing out fibre as is done in equation 42. However, with higher order coincidence that account for the summing in there is a degeneracy in the basis vectors for different detection scenarios. For example, the basis vector $\widehat{p_{32}}\otimes\widehat{p_{21}}\otimes\widehat{p_{10}}$ could correspond to the summing-in to $\widehat{p}_{31}$ in coincidence with $\widehat{p}_{10}$ or $\widehat{p}_{32}$ in coincidence with the summing-in to $\widehat{p}_{20}$. Calculating the coincidence vector as shown in equation 49 would provide the probability for both of these distinct detection scenarios onto the same basis vector. To avoid this detection degeneracy ( distinct detection events possessing the same basis vector), a gate projector can be applied to the gated hit detection map vector after power expansion.
IV Remarks
IV.1 Auxiliary Radiation
A note should be said here on what motivated this framework. As previously mentioned, the transition matrix formalism for summing corrections in nuclear level schemes was initially put together by Semkow. Afterwards, others like [5] extended it by including auxiliary radiation; that is, others forms of radiation like X-rays from internal conversion and 511 keV gamma rays from position-electron annihilation. It is typical that auxiliary radiation can be added on into the formalism by the use of virtual branches. In this case, branches are added on to the decay level scheme with the additional transitions representing the auxiliary radiation. There are multiple ways to do this corresponding to different types of radiation. For X-rays, [5] showed that virtual branches can be placed between the levels to allow for internal conversion and X-rays. In,[6]–predating the work of [8]– 511 keV gammas are proposed to be incorporated as ”pseudo-branches” above the level scheme, where internal conversion coefficients of $0.5$ double the probability of the emission allowing for the two 511 keV produced in annihilation–though not to be detected simultaneously given a single detector setup. For modern day gamma spectroscopy experiments, like those done with [3], a spherical distribution of multiple detectors is more standard and allows for coincidence measurements. For multiple detector arrangements having two 511 keV gammas below the ground state, as shown in Figure 3, is more suitable. In this manner, we ensure that all decays give the 511 keV radiation from the positron annihilation. The transition matrix formalism can handle the summing out corrections with the 511 virtual branches but is restricted to only allow summing in corrections with the transitions directly connected above the ground state. It is this very need to allow for full energy summing in corrections with 511 keV gammas in general that has motivated this work. An example where this is important is the $\beta^{+}$ decay of $^{22}Mg$ where summing with 511 keV gammas corrupts one of the large peaks of interest  [^3] The precision measurement of the superallowed $\beta^{+}$ decay of $^{22}Mg$–a measurement that contributes to the $V_{ud}$ value in the CKM matrix–requires a precise measurement of this corrupted peak.. However, the coincidence algebra not only addresses this by allowing for full energy summing-in between all gammas, but further separates coincidence terms by their branches and makes the extension to gated probabilities trivial.
Figure 3: A general nuclear level scheme with virtual branches below the ground state to allow for 511 keV gammas in the summing formalism.
IV.2 Conclusion
@@
Motivated by the need to extend the generality of the transition matrix formalism traditionally used to model nuclear decay schemes, We have demonstrated how the path algebra of a decay quiver gives the underlying algebra for transitions matrices. This allowed for an extension into a tensor decay space which we have restricted by defining a coincidence vector space upon which we build the coincidence algebra. The coincidence algebra allows for a more general algebra for calculating coincidence probabilities as it allows for multiplication between non connected paths in the decay quiver and in addition separates different branching terms. With the path algebra being the space with which nuclear levels schemes live, we have define an algebra bundle over the path algebra where each decay vector defines the structure constants on the coincidence algebra at the fibre. We have showed, at the very least that this algebra allows for additional coincidence probability terms to be calculated. We propose, however, that this method of taking fibre bundles of quivers, or potentially even quiver representations, may be a fruitful method for modeling all nuclear levels schemes.